OpenStudy (anonymous):

Please HELP!! Is pi/4 a solution for 3 sqrt 2 cos theta + 2 = -1?

OpenStudy (anonymous):

\[3\sqrt{2\cos \theta+2} = -1\] does it look like this

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

cos theta = -19/81 so pi/4 is not the answer

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem